One base of the trapezoid is 1 ft shorter than the other base. Find the lengths of teh bases if the trapezoid is 4 ft high and has an area of 10 ft sq

Answer by Boreal(15235)

A trap=(1/2)h (b1+b2)
b1=b2-1
10=(1/2)*4(b2-1+b2)
5=2b2-1
2b2=6
b2=3 ft ANSWER
b1=2 ft ANSWER
A=(1/2)(5)(4)=10 sq ft